OpenStudy (anonymous):

3SinA-4CosA=0; Find SinA. How to simplify without the angles? help!

OpenStudy (anonymous):

3SinA-4CosA=0 (3SinA-4CosA)(3SinA+4CosA)=0 9Sin^2 A-16Cos^2 A=0 9Sin^2 A-16(1-Sin^2 A)=0 25Sin^2 A = 16 Sin^2 A = 16/25 Sin A = 4/5 or -4/5 (rejected)

OpenStudy (anonymous):

how come this step? (3SinA-4CosA)(3SinA+4CosA)=0

OpenStudy (anonymous):

(3SinA-4CosA)(3SinA+4CosA)=0 * (3SinA+4CosA)

OpenStudy (anonymous):

Oh Thanks!! :D

OpenStudy (anonymous):

@Davidc